Please show work. Thanks.


4. Given an effective annual interest rate of 12%, what is the nominal interest rate compounded quarterly? A) 11.5 B) 12.0 C) 12.4 D 12.5
Answer #1

Solution:

Let nominal rate of interest is = 4i per annum

Quarterly rate = 4i/4 = i

Lets principal amount is $100

Future value after 1 year using effective rate = $100*112% = $112

Now

$100 (1+i)^4 = $112

(1+i)^4 = 1.12

i = 2.875%

Nominal rate of interest = 2.875*4 = 11.50%

Hence option A is correct.